OpenStudy (anonymous):

complete the square 2x2-7x=-3

OpenStudy (anonymous):

4-7x=-3 -4 -4 7x= -7 x= -1

hartnn (hartnn):

its 2x^2 not 2*x*2

OpenStudy (anonymous):

? .... hu?

hartnn (hartnn):

\(2x^2\)

hartnn (hartnn):

\(2x^2-7x=-3\)

OpenStudy (anonymous):

Ohh

hartnn (hartnn):

now can u help sam ?

OpenStudy (anonymous):

I dont think so, its hard & im kind of busy, i just wanted to see if i was any help. Can you help sam?

OpenStudy (anonymous):

meh ill do it ALRIGHT , first lets remove that 2 infront of the x^2 divide the entire thing by 2 call me when you've done that (and show me the resulting equation)